OpenStudy (anonymous):

last question!!! (: the company discovered that it costs $30 to produce 2 widgets, $106 to produce 4 widgets, and $670 to produce 10 widgets. Using the quadratic function, find the cost of producing 5 widgets.

OpenStudy (anonymous):

cost = ax^2+bx+c 30 = 4a+2b+c 106 = 16a +4b +c 670 = 100a +10b +c solve the equations to get a,b,c

OpenStudy (amistre64):

dhats got a good start for us...

OpenStudy (amistre64):

-30 = - 4a -2b -c -30 = -4a -2b -c 106 = 16a +4b +c 670 = 100a +10b +c ---------------------------------------------- 76 = 12a +2b 640 = 96a +8b

OpenStudy (amistre64):

640 = 96a +8b -608 = -96a -16b ------------------ 32 = -8b b = -4

OpenStudy (anonymous):

i got a = 7

OpenStudy (amistre64):

a = 76-2b//12 a = 84/12 = 7 good job....now lets hope its right :)

OpenStudy (amistre64):

30 = 4(7)+2(-4)+c

OpenStudy (anonymous):

One thing I don't get about these questions is why buying more in total increases the cost per thing? As far as I recall, all have been like this ... :(

OpenStudy (amistre64):

30 = 28 -8 +c 30 = 20 + c 10 = c 7x^2 -4x +10 = y should work, lets check ourselves

OpenStudy (amistre64):

economy of scale :)

OpenStudy (anonymous):

Economy of scale... in reverse, though, more like :/

OpenStudy (amistre64):

2,30 4,106 10,670 700 -40 +10 = 700 - 30 = 670..checks 7(16) -4(4) + 10 = 112-16+10 = 112 - 6 = 106 checks 28 -8 + 10 = 28+2 = 30 its good

OpenStudy (amistre64):

7(5)^2 -4(5) +10 = y 7(25) -20 +10 = y 175 -10 = 165 should be our answer...

OpenStudy (amistre64):

yeah....in reverse :)

OpenStudy (anonymous):

THANK YOU SO MUCH!!! do you have an email i could get ahold of you on? just in case i need help in other stuff.

OpenStudy (amistre64):

lol....youre welcome, but I think ill just stick to openstudy stuff ;)

OpenStudy (amistre64):

maybe widgets are one of those products that consumes more than its worth....like tribbles :)

OpenStudy (anonymous):

¬_¬

OpenStudy (anonymous):

A Mathematica approach: let data = {{2, 30}, {4, 106}, {10, 670}} Fit[ data, {1, x, x^2}, x] /. x -> 5 yeilds 10. - 4. x + 7. x^2 =165 at x=5 The Fit function is a curve fitter.

OpenStudy (amistre64):

i never trust a machine to do a "man"s job. unless its a slide rule :)

OpenStudy (anonymous):

Why the quotation marks around man?

OpenStudy (amistre64):

trying to preserve my anonymity lol

OpenStudy (anonymous):

:(

OpenStudy (amistre64):

i try to remain androgenous online ;)

OpenStudy (anonymous):

As a 76 year old person who could not integrate x^2 with pencil and paper I am very impressed with amistre64 solving this thing in real time.

OpenStudy (anonymous):

Hmm, but which are you ¬_¬

OpenStudy (anonymous):

You are 76?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

But you can integrate x^2, right?

OpenStudy (amistre64):

if i had done this with pencil and paper i would have been done long time ago.... but trying to remember which keys to press is a challange

OpenStudy (anonymous):

I graduated in the lower 1/3 of my class at San Jose State College as a Math major. I cannot remember all of the rules. Rather I'm using software developed by a group of World class mathematicians and computer programmers. Using Mathematic 8. The installation "file" is about 3Gb in size. By the way, Wolfram practically gives their system to students and acedemics.

OpenStudy (anonymous):

Hmm, intriguing. I personally find it easier to learn the rules than to learn what format to type things into Wolfram|Alpha (though maybe Mathematica is slightly easier to use). Can of course be useful at times, though.

OpenStudy (amistre64):

i just make up my own rules half the time, and guess at the the rest :)

OpenStudy (anonymous):

Got to go now. Thanks for the exchange.